OpenStudy (anonymous):

What is the seventh term of the sequence 1, 3, 9, 27, ...? 81 243 729 2187

OpenStudy (the_fizicx99):

What is 27/9 ?

OpenStudy (anonymous):

3 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Yes, lol So what is 9/3 ?

OpenStudy (anonymous):

1 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

9/3 is 1 ??! O_O

OpenStudy (anonymous):

yes?? @tHe_FiZiCx99

OpenStudy (mathmale):

9 DIVIDED BY 3 IS 1??

OpenStudy (anonymous):

-3??

OpenStudy (mathmale):

Perhaps someone intended that to be a joke and I was too slow to get it. :)

OpenStudy (anonymous):

I got 1

OpenStudy (anonymous):

81

OpenStudy (the_fizicx99):

No, its 3 and 3/1 is 3 xD so their common ratio is 3 So easily plug in for: f(n) = f(1) • r^n-1, Plug in any number you want to find, I'll find the 6th term. f(6) = f(1) • 3^6-1 f(6) = f(1) • 3^5 f(6) = f(1) • 243 3 • 243 f(6) = 729

OpenStudy (anonymous):

a(n) = 3^(n-1) for all real numbers n

OpenStudy (the_fizicx99):

My bad, f(1) = 1 1 * 243 = 243 f(6) = 243 f(7) = 729 lol now there's your answer :#

OpenStudy (mathmale):

larryboxaplenty: Great that you could come up with such an appropriate formula!! As for n values, though, wouldn't the domain be "integers starting with 1", and not "all real n"?